OpenStudy (dannyboi):

Find the common difference of the arithmetic sequence shown. 1/4, 3/8, 1/2 1) 1/16 2)1/8 3)1/4

OpenStudy (jango_in_dtown):

subtract the first number from the second number 3/8-1/4=(3-2)/8=1/8

OpenStudy (jango_in_dtown):

similary subtract the second number from the third number

OpenStudy (dannyboi):

ok

OpenStudy (jango_in_dtown):

see it will be same.. it will be the common difference

OpenStudy (dannyboi):

ok thanx